OpenStudy (anonymous):

A bullet of mass 45 g is fired at a speed of 220 m/s into a 5.0-kg sandbag hanging from a string from the ceiling. The sandbag absorbs the bullet and begins to swing. To what maximum vertical height will it rise? Hint: Energy is not conserved as the bullet enters the bag but is conserved after the bullet comes to rest in the bag and the bag is swinging upward

OpenStudy (anonymous):

(m1*V1)=(m1+m2)Vnew Vnew=1.96m/s After that I'm stuck

OpenStudy (anonymous):

You may use the principle of censervation of energy to find the height. \[E=E _{p}+E _{k}=0\] \[E _{p}=m \times g \times h\] \[E _{k}=\frac{ 1 }{ 2 } \times m \times v ^{2}\] \[E _{p}=-E _{k}\] replacing the data in the formulas, you might obtain the height of the sandbag.